OpenStudy (anonymous):

Solve P = 2(l + w) for l. What are the missing values in the table? (1 point) P w l 14 2 ? 22 8 ? Create your own unique literal equation and solve for one of the variables. Show your work. Then, using complete sentences, explain how you solved for the variable you chose. (2 points) Using complete sentences, explain how solving a literal equation is similar to or different from simplifying an expression such as 6 - 2(52 + 7) � 4. (2 points) Using complete sentences, explain what might happen if the order of operations was used to solve a literal equation. (1 point)

OpenStudy (anonymous):

help please!!!

OpenStudy (anonymous):

To solve for l in terms of the other literals, rewrite the given literal equation, using the distributive law of multiplication over addition: p = 2l + 2w. Py the wymmetry property of equality we may rewrite it as 2l + 2w = p. Subtracting 2w from both sides to isolate the 2l results in 2l = p - 2w. Dividing both sides by 2 isolates i: l = p/2 - w, and we have solved for l in terms of the other literals.

OpenStudy (anonymous):

From the second row in the table we have p = 12 and w = 2. Substituting those numerical values into the last literl equation gives l = 14/2 - 2, which evaluates to 5. That answers the first question in the table. A similar calculation using the imformation in the third row will show that l = 3, answering the second question in the table.